2- What is the pH for a 0.0140 M solution of sodium acetate (NaAc). The pKa value for acetic acid is 4.75. ANS: ...
Answer #1

Solution-
The reaction can be written as

CH3COONa + H2O -------> CH3COOH + OH- + Na+

0.0140 - X X X

Now the pKa for CH3COONa = 4.75

So the Kb = 10^-pKb = 1.78*10^-5

=>Kb = 1.78*10^-5 = X^2 / 0.0140 - X
=>1.78*10^-5*0.0140 + 1.78*10^-5*(-x)= x^2

=> X^2+1.78*10^-5x -2.492e-7=0

=> X = 0.000490 M = [OH-]

pOH = - log [OH-] = - log (0.000490)

= 3.30

pH = 14 - 3.30

= 10.7